# Gaussian-Based Periodic Grand Canonical Density Functional Theory with Implicit Solvation for Computational Electrochemistry

### v2

@@ Abstract
We present a numerical method for grand canonical density functional theory (DFT) tailored to solid-state systems, employing Gaussian-type orbitals as the primary basis. Our approach directly minimizes the grand canonical free energy using the density matrix as the sole variational parameter, while self-consistently updating the electron number between self-consistent field iterations. To enable realistic electrochemical modeling, we integrate this approach with implicit solvation models. Our solvation scheme introduces less than 50% overhead relative to gas-phase calculations. Compared to existing plane wave-based implementations, our method shows improved robustness in grand canonical simulations. We validate the approach by modeling corrosion at silver surfaces, finding excellent agreement with previous studies. This work lays the groundwork for future wavefunction-based simulations beyond DFT under electrochemical operando conditions.
@@ I Introduction
Electrochemical reactions at solid-liquid interfaces are the cornerstone of many critical processes for renewable energy, such as carbon dioxide reduction and water splitting.Walter et al. (2011); Seh et al. (2017) As a result, there has been a considerable interest in designing more efficient electrochemical cells and catalysts to perform these reactions.Gür (2018) In parallel with experimental progress, there is a growing demand for accurate computational modeling of the chemistry and physics occurring at solid-liquid interfaces.Nørskov et al. (2011); Chen et al. (2019); Hammes-Schiffer and Galli (2021); Ringe et al. (2022) Because interfacial electrochemical reactions inherently involve electron transfer—often accompanied by bond formation and breaking—accurately modeling these processes requires a quantum mechanical description of the electrons at the interface.
@@
One challenge of quantum mechanically modeling solid-liquid interfaces is that an atomistic model is computationally infeasible even with a relatively economical method such as density functional theory (DFT).Bruix et al. (2019) Fortunately, chemical reactivity in these systems is often localized to the environment near the interface. Therefore, only the portion of the electrode close to the surface needs to be explicitly modeled. To maintain computational tractability, the solvent and electrolyte environment can be coarse-grained down to a polarizable continuum with an implicit solvation model.Ringe et al. (2022) These methods in general work by defining a solvent-accessible cavity that influences the electrostatic potential in the solute environment. Ionic screening is also included to account for the influence of electrolyte ions in solution. These methods have been implemented in a wide range of periodic quantum chemistry codes based on plane waves.Mathew et al. (2019); Fisicaro et al. (2016); Letchworth-Weaver and Arias (2012); Dziedzic et al. (2020)
@@
Moreover, electrochemical reactions are often performed under a constant electrochemical potential. Under these conditions, one aims to minimize the grand canonical free energy variationally, as opposed to the minimization of the electronic energy. This is challenging as most electronic structure methods available to date are formulated for canonical ensembles and assume constant electron number. In some past works, a “constant potential" calculation was performed by running multiple fixed-electron-number calculations and interpolating to obtain the proper chemical potential.Bonnet et al. (2012); Gauthier et al. (2019) These methods require additional computational or manual effort, so a technique that avoids these extra complications is highly desirable. For this reason, previously, direct minimization methods with an auxiliary Hamiltonian and related methods have been developed in plane wave density functional theory codes.Marzari, Vanderbilt, and Payne (1997); Freysoldt, Boeck, and Neugebauer (2009); Sundararaman, Goddard, and Arias (2017); Melander et al. (2018) We also note a few exceptions for many-body methods such as perturbation theory,Hirata and Jha (2020); Wei et al. (2022) coupled-cluster methods,Baeurle, Martoňák, and Parrinello (2002); Hummel (2018); White and Chan (2018); White and Kin-Lic Chan (2020); Harsha, Henderson, and Scuseria (2019) quantum Monte Carlo,Shen et al. (2020); Malone et al. (2015); Lee, Morales, and Malone (2021) etc., where grand canonical electronic structure methods were developed focusing on studying warm dense matter in the gas phase.
@@
Traditionally, periodic electronic structure calculations are performed with plane waves as the basis of choice.Kresse and Furthmüller (1996); Giannozzi et al. (2009); Sundararaman et al. (2017) This is because plane waves are inherently periodic and form an orthogonal basis set, making them free from basis set superposition error (BSSE). On the other hand, Gaussian-type orbitals (GTOs), the natural choice for molecular applications, are non-orthogonal and local. They offer a more compact representation than plane waves for performing many-body calculations. Therefore, despite their deficiencies, such as BSSE and basis set linear dependencies,Klahn and Bingel (1977); Vilela Oliveira et al. (2019) GTOs offer unique strengths, especially beyond pure density functional calculations.Pisani et al. (2012, 2008); Rybkin and VandeVondele (2016); Robinson et al. (2025) Recently, work has been done to develop GTO basis sets that extrapolate well to the basis set limit, making GTOs a much more viable choice as the computational basis in solids.Daga, Civalleri, and Maschio (2020); Peintinger, Oliveira, and Bredow (2013); Ye and Berkelbach (2022)
@@
This paper is organized as follows. We first present our implementation of periodic grand canonical density functional theory (GCDFT) in a GTO-based code that adaptively updates the electron number between self-consistent field (SCF) iterations. Our method performs a variational minimization of the grand potential over density matrices, as initially proposed by Mermin in the celebrated thermal Hartree-Fock theory paper.Mermin (1963) Such an implementation is impractical in plane wave frameworks due to the significant storage overhead associated with density matrices. Furthermore, our method does not require simultaneous optimization of both molecular orbital coefficients and an auxiliary Hamiltonian used in plane wave implementations, which, as we demonstrate later, shows faster convergence than preexisting plane wave methods.Marzari, Vanderbilt, and Payne (1997); Freysoldt, Boeck, and Neugebauer (2009); Sundararaman, Goddard, and Arias (2017)
@@
We also present the implementation of two implicit solvation models, which incorporate linear dielectric and ionic response into the density functional. Both models have demonstrated accuracy on a wide range of solutes and are seamlessly integrated into our GCDFT implementation.Gunceler et al. (2013); Sundararaman and Goddard (2015) We note that while CP2K has an implementation of grand canonical density functional theory with the self-consistent continuum solvation (SCCS) model,Andreussi, Dabo, and Marzari (2012) they employ a Pulay mixing scheme rather than a more robust variational minimization method.Chai and Luber (2024) To the best of our knowledge, no existing GTO-based implementation supports variational minimization of the grand canonical free energy in the presence of solvation under periodic boundary conditions. Our resulting GCDFT method enables realistic simulations of electrochemical systems. We present an application of our method to a detailed study of the corrosion of a silver electrode surface.Kang et al. (2024) This work demonstrates the viability, robustness, and potential of using GTOs as a computational basis for practical problems in electrochemistry, which is currently dominated by plane wave-based codes.Goodpaster, Bell, and Head-Gordon (2016); Garza, Bell, and Head-Gordon (2018)
@@ II.1 Grand Canonical Density Functional Theory
We discretize the underlying operators in Eq.˜1 with orthonormalized GTOs,
$\frac{\Omega}{N_{k}}=\frac{1}{N_{k}}\sum_{\mathbf{k}}\left(\frac{1}{2}\text{Tr}({\textbf{P}^{\mathbf{k}}}({\textbf{h}^{\mathbf{k}}}+{\textbf{F}^{\mathbf{k}}}))-\mu\text{Tr}({\textbf{P}}^{\mathbf{k}})+\frac{1}{\beta}S_{\text{el}}^{\mathbf{k}}\right),$ (2)
where $\mathbf{k}$} is a $\mathbf{k}$-point in the Brillouin zone, $N_{k}$ is the number of $\mathbf{k}$-points, ${\textbf{P}}^{\mathbf{k}}$ is the one-body reduced density matrix (1RDM) at $\mathbf{k}$, ${\textbf{h}}^{\mathbf{k}}$ and ${\textbf{F}}^{\mathbf{k}}$ are the one-body Hamiltonian matrix and the Fock matrix at $\mathbf{k}$, respectively, and $S_{\text{el}}^{\mathbf{k}}=\text{Tr}({\textbf{P}}^{\mathbf{k}}\log({\textbf{P}}^{\mathbf{k}})+(\textbf{I}-{\textbf{P}}^{\mathbf{k}})\log(\textbf{I}-{\textbf{P}}^{\mathbf{k}}))$. In practice, we use the transformation between non-orthogonal atomic orbitals and orthogonalized atomic orbitals to correctly handle the metric when evaluating Eq.˜2 with ${\textbf{h}}^{\mathbf{k}}$ and ${\textbf{F}}^{\mathbf{k}}$ evaluated in the atomic orbital basis.
@@
${\textbf{P}}^{\mathbf{k}}=\frac{1}{\exp(\text{\boldmath$\Gamma$}^{\mathbf{k}})+1}=\textbf{U}^{\mathbf{k}}\frac{1}{\exp(\text{\boldmath$\gamma$}^{\mathbf{k}})+1}\textbf{U}^{\mathbf{k}H}=\textbf{U}^{\mathbf{k}}\text{\boldmath$\rho$}^{\mathbf{k}}\textbf{U}^{\mathbf{k}H},$ (3)
where $\text{\boldmath$\Gamma$}^{\mathbf{k}}=\beta\left(\frac{1}{2}\left({\textbf{H}}^{\mathbf{k}}+{\textbf{H}}^{\mathbf{k}H}\right)-\mu\textbf{I}\right)=\textbf{U}^{\mathbf{k}}\text{\boldmath$\gamma$}^{\mathbf{k}}\textbf{U}^{\mathbf{k}H}$. We can then minimize the free energy expression by performing an unconstrained minimization with respect to $\{{\textbf{H}}^{\mathbf{k}}\}$. At convergence, we have ${\textbf{H}}^{\mathbf{k}}={\textbf{F}}^{\mathbf{k}}$. We note that this method differs from the popular auxiliary Hamiltonian approaches in that we do not optimize the coefficient matrix as a separate variational parameter.Marzari, Vanderbilt, and Payne (1997); Freysoldt, Boeck, and Neugebauer (2009); Sundararaman, Goddard, and Arias (2017) Instead, our molecular orbital coefficient matrix, $\{\mathbf{C}^{\mathbf{k}}\}$, is determined by diagonalizing $\text{\boldmath$\Gamma$}^{\mathbf{k}}$. Additionally, due to the compactness of GTOs, we can work with the full density matrix rather than a subspace of the given basis set as is done in plane wave-based codes. This is because in plane wave-based codes, the number of basis functions is typically so large that it becomes infeasible to store the full density matrix. While a recent effort was made to perform direct minimization in the presence of fluctuating orbital population, the method is applicable only to systems with integer electrons.Li et al. (2025)
We perform the minimization of the free energy using the nonlinear conjugate gradient method. In the method, we use the residual $\frac{1}{2}({\textbf{F}}^{\mathbf{k}}-{\textbf{H}}^{\mathbf{k}})$ as the preconditioned gradient.Sundararaman, Goddard, and Arias (2017) We note that this method converges to the true minimum since ${\textbf{F}}^{\mathbf{k}}-{\textbf{H}}^{\mathbf{k}}=0$ at convergence. The gradient of the objective function $\Omega$ with respect to $\text{\boldmath$\Gamma$}^{\mathbf{k}}$ can be computed efficiently, and the gradient with respect to ${\textbf{H}}^{\mathbf{k}}$ is obtained by symmetrizing the gradients with respect to $\text{\boldmath$\Gamma$}^{\mathbf{k}}$. The detailed forms of the gradients are included below. All expressions are written using the Einstein summation convention. We define $G_{li}^{\textbf{k}}:=\frac{1}{\epsilon_{i}^{\textbf{k}}-\epsilon_{l}^{\textbf{k}}}$ where $\epsilon_{i}^{\textbf{k}}$ is the $i$-th eigenvalue of the Fock matrix, $\mathbf{F}^{\mathbf{k}}$.
@@
To get the gradient with respect to ${\textbf{H}}^{\textbf{k}}$ from the gradient with respect to $\text{\boldmath$\Gamma$}^{\textbf{k}}$, we can use the following:
$\frac{\partial\Omega}{\partial\text{Re}({\textbf{H}}^{\textbf{k}})}=\frac{\beta}{2}\left(\frac{\partial\Omega}{\partial\text{Re}\left(\text{\boldmath$\Gamma$}^{\textbf{k}}\right)}+\left(\frac{\partial\Omega}{\partial\text{Re}\left(\text{\boldmath$\Gamma$}^{\textbf{k}}\right)}\right)^{T}\right)$ (6)
$\frac{\partial\Omega}{\partial\text{Im}({\textbf{H}}^{\textbf{k}})}=\frac{\beta}{2}\left(\frac{\partial\Omega}{\partial\text{Im}\left(\text{\boldmath$\Gamma$}^{\textbf{k}}\right)}-\left(\frac{\partial\Omega}{\partial\text{Im}\left(\text{\boldmath$\Gamma$}^{\textbf{k}}\right)}\right)^{T}\right)$ (7)
With this method, we can converge systems at a wide range of chemical potentials.
Note that this method can also be used to perform fixed-electron calculations as in canonical density functional theory by adding the gradient of the chemical potential $\mu$ with respect to the variational parameters and using the canonical free energy $E=E_{\text{DFT}}-TS_{\text{el}}$ as the objective function for optimization.Freysoldt, Boeck, and Neugebauer (2009) This algorithm was also used to obtain the potential of zero charge (PZC) of metal slabs later in this work. The relevant gradients are as follows:
@@ II.2 Implicit Solvation Framework
Meaningful constant potential calculations in periodic systems must resolve one problem inherent to periodic boundary conditions. That is, the potential energy of charged unit cells in an infinite lattice diverges. In nearly all periodic electronic structure codes, the default way to handle this is by zeroing out the $G=0$ component of the electrostatic potential in reciprocal space, which is equivalent to introducing a uniform background charge that neutralizes the charge of the system.Giannozzi et al. (2009); Kresse and Furthmüller (1996) It is known that the choice of neutralizing charge can affect qualitative predictions, so a physically informed choice should be used in calculations.Bhandari et al. (2020) As electrochemical reactions are performed in the presence of an electrolyte solution, the most logical way to enforce the net charge neutrality is by introducing ionic screening provided by electrolytes in solution. It is well-known that introducing linear ionic response (i.e., Debye screening) is already sufficient to neutralize the unit cell exactly.Letchworth-Weaver and Arias (2012)
In this work, we have implemented a density-based polarizable continuum solvation model (PCM), which incorporates linear dielectric and ionic response. We demonstrate the implementation of two cavity functions. The first is a locally determined cavity based on the local linear PCM model (LPCM) in JDFTx and VASPsol.Gunceler et al. (2013); Mathew et al. (2019) The LPCM model is equivalent to computing the solvation potential via the linearized Poisson-Boltzmann equation. Secondly, we have implemented the charge-asymmetric nonlocally-determined local-electric (CANDLE) cavity function, known for its prediction of accurate energies for charged solutes, especially anions, which are systematically undersolvated in most PCM models.Sundararaman and Goddard (2015) We provide more implementation details of these solvation models below.

### v3

@@ Abstract
We present a numerical method for grand canonical density functional theory (DFT) tailored to solid-state systems, employing Gaussian-type orbitals as the primary basis. Our approach directly minimizes the grand canonical free energy using the density matrix as the sole variational parameter, while self-consistently updating the electron number between self-consistent field iterations. To enable realistic electrochemical modeling, we integrate this approach with implicit solvation models. Our solvation scheme introduces less than 50% overhead relative to gas-phase calculations. Compared to existing plane wave-based implementations, our method shows improved robustness in grand canonical simulations. We validate the approach by modeling corrosion at silver surfaces, finding excellent agreement with previous studies. Our method is implemented in the quantum chemistry software Q–Chem. This work lays the groundwork for future wavefunction-based simulations beyond DFT under electrochemical operando conditions.
@@ I Introduction
Electrochemical reactions at solid-liquid interfaces are the cornerstone of many critical processes for renewable energy, such as carbon dioxide reduction and water splitting.1, 2 As a result, there has been a considerable interest in designing more efficient electrochemical cells and catalysts to perform these reactions.3 In parallel with experimental progress, there is a growing demand for accurate computational modeling of the chemistry and physics occurring at solid-liquid interfaces.4, 5, 6, 7 Because interfacial electrochemical reactions inherently involve electron transfer—often accompanied by bond formation and breaking—accurately modeling these processes requires a quantum mechanical description of the electrons at the interface.
@@
One challenge of quantum mechanically modeling solid-liquid interfaces is that an atomistic model is computationally infeasible even with a relatively economical method such as density functional theory (DFT).8 Fortunately, chemical reactivity in these systems is often localized to the environment near the interface. Therefore, only the portion of the electrode close to the surface needs to be explicitly modeled. To maintain computational tractability, the solvent and electrolyte environment can be coarse-grained down to a polarizable continuum with an implicit solvation model.7 These methods in general work by defining a solvent-accessible cavity that influences the electrostatic potential in the solute environment. Ionic screening is also included to account for the influence of electrolyte ions in solution. These methods have been implemented in a wide range of periodic quantum chemistry codes based on plane waves.9, 10, 11, 12
@@
Moreover, electrochemical reactions are often performed under a constant electrochemical potential. Under these conditions, one aims to minimize the grand canonical free energy variationally, as opposed to the minimization of the electronic energy. This is challenging as most electronic structure methods available to date are formulated for canonical ensembles and assume constant electron number. In some past works, a “constant potential” calculation was performed by running multiple fixed-electron-number calculations and interpolating to obtain the proper chemical potential.13, 14 These methods require additional computational or manual effort, so a technique that avoids these extra complications is highly desirable. For this reason, previously, direct minimization methods with an auxiliary Hamiltonian and related methods have been developed in plane wave density functional theory codes.15, 16, 17, 18 We also note a few exceptions for many-body methods such as perturbation theory,19, 20 coupled-cluster methods,21, 22, 23, 24, 25 quantum Monte Carlo,26, 27, 28 etc., where grand canonical electronic structure methods were developed focusing on studying warm dense matter in the gas phase.
@@
Traditionally, periodic electronic structure calculations are performed with plane waves as the basis of choice.29, 30, 31 This is because plane waves are inherently periodic and form an orthogonal basis set, making them free from basis set superposition error (BSSE). On the other hand, Gaussian-type orbitals (GTOs), the natural choice for molecular applications, are non-orthogonal and local. They offer a more compact representation than plane waves for performing many-body calculations. Therefore, despite their deficiencies, such as BSSE and basis set linear dependencies,32, 33 GTOs offer unique strengths, especially beyond pure density functional calculations.34, 35, 36, 37 Recently, work has been done to develop GTO basis sets that extrapolate well to the basis set limit, making GTOs a much more viable choice as the computational basis in solids.38, 39, 40
@@
This paper is organized as follows. We first present our implementation of periodic grand canonical density functional theory (GCDFT) in a GTO-based code that adaptively updates the electron number between self-consistent field (SCF) iterations. Our method performs a variational minimization of the grand potential over density matrices, as initially proposed by Mermin in the celebrated thermal Hartree-Fock theory paper.41 Such a direct implementation is impractical in plane wave frameworks due to the significant storage overhead associated with density matrices. Furthermore, our method does not require simultaneous optimization of both molecular orbital coefficients and an auxiliary Hamiltonian used in plane wave implementations, which, as we demonstrate later, shows faster convergence than preexisting plane wave methods.15, 16, 17
@@
We also present the implementation of two implicit solvation models, which incorporate linear dielectric and ionic response into the density functional. Both models have demonstrated accuracy on a wide range of solutes and are seamlessly integrated into our GCDFT implementation.42, 43 We note that while CP2K has an implementation of grand canonical density functional theory with the self-consistent continuum solvation (SCCS) model,44 they employ a Pulay mixing scheme rather than a more robust variational minimization method.45, 46 The CP2K implementation uses a planar counter charge formalism for ionic screening rather than a more realistic implicit electrolyte. To the best of our knowledge, no existing GTO-based implementation supports variational minimization of the grand canonical free energy in the presence of solvation under periodic boundary conditions. Our resulting GCDFT method enables realistic simulations of electrochemical systems. We present an application of our method to a detailed study of the corrosion of a silver electrode surface.47 This work demonstrates the viability, robustness, and potential of using GTOs as a computational basis for practical problems in electrochemistry, which is currently dominated by plane wave-based codes.48, 49
@@ II.1 Grand Canonical Density Functional Theory
We discretize the underlying operators in Eq. 1 with orthonormalized GTOs,
$\frac{\Omega}{N_{k}}=\frac{1}{N_{k}}\sum_{\mathbf{k}}\left(\text{Tr}({\textbf{P}^{\mathbf{k}}}({\textbf{h}^{\mathbf{k}}}+\frac{1}{2}{\textbf{J}^{\mathbf{k}}}))+E_{\text{xc}}[\rho(\mathbf{r}),\cdots]-\mu\text{Tr}({\textbf{P}}^{\mathbf{k}})+\frac{1}{\beta}S_{\text{el}}^{\mathbf{k}}\right),$ (2)
where $\mathbf{k}$} is a $\mathbf{k}$-point in the Brillouin zone, $N_{k}$ is the number of $\mathbf{k}$-points, ${\textbf{P}}^{\mathbf{k}}$ is the one-body reduced density matrix (1RDM) at $\mathbf{k}$, ${\textbf{h}}^{\mathbf{k}}$ and ${\textbf{J}}^{\mathbf{k}}$ are the one-body Hamiltonian matrix and the Coulomb matrix at $\mathbf{k}$, respectively, $E_{\text{xc}}$ is the exchange-correlation energy, and $S_{\text{el}}^{\mathbf{k}}=\text{Tr}({\textbf{P}}^{\mathbf{k}}\log({\textbf{P}}^{\mathbf{k}})+(\textbf{I}-{\textbf{P}}^{\mathbf{k}})\log(\textbf{I}-{\textbf{P}}^{\mathbf{k}}))$. In practice, we use the transformation between non-orthogonal atomic orbitals and orthogonalized atomic orbitals to correctly handle the metric when evaluating Eq. 2 with ${\textbf{h}}^{\mathbf{k}}$ and ${\textbf{F}}^{\mathbf{k}}$ evaluated in the atomic orbital basis.
@@
${\textbf{P}}^{\mathbf{k}}=\frac{1}{\exp(\text{$\Gamma$}^{\mathbf{k}})+1}=\textbf{U}^{\mathbf{k}}\frac{1}{\exp(\text{$\gamma$}^{\mathbf{k}})+1}\textbf{U}^{\mathbf{k}H}=\textbf{U}^{\mathbf{k}}\text{$\rho$}^{\mathbf{k}}\textbf{U}^{\mathbf{k}H},$ (3)
where $\text{$\Gamma$}^{\mathbf{k}}=\beta\left(\frac{1}{2}\left({\textbf{H}}^{\mathbf{k}}+{\textbf{H}}^{\mathbf{k}H}\right)-\mu\textbf{I}\right)=\textbf{U}^{\mathbf{k}}\text{$\gamma$}^{\mathbf{k}}\textbf{U}^{\mathbf{k}H}$. We can then minimize the free energy expression by performing an unconstrained minimization with respect to $\{{\textbf{H}}^{\mathbf{k}}\}$. At convergence, we have ${\textbf{H}}^{\mathbf{k}}={\textbf{F}}^{\mathbf{k}}$. We note that this method differs from the popular auxiliary Hamiltonian approaches in that we do not optimize the coefficient matrix as a separate variational parameter.15, 16, 17 Instead, our molecular orbital coefficient matrix, $\{\mathbf{C}^{\mathbf{k}}\}$, is determined by diagonalizing $\text{$\Gamma$}^{\mathbf{k}}$. Additionally, due to the compactness of GTOs, we can work with the full density matrix rather than a subspace of the given basis set as is done in plane wave-based codes. This is because in plane wave-based codes, the number of basis functions is typically so large that it becomes infeasible to store the full density matrix. Our method incurs minimal storage overhead, as {H} has the same dimension as the density matrix. While a recent effort was made to perform direct minimization in the presence of fluctuating orbital population, the method is applicable only to systems with integer electrons.50
We perform the minimization of the free energy using the nonlinear conjugate gradient method. In the method, we use the residual $\frac{1}{2}({\textbf{F}}^{\mathbf{k}}-{\textbf{H}}^{\mathbf{k}})$ as the preconditioned gradient.17 We note that this method converges to the true minimum since ${\textbf{F}}^{\mathbf{k}}-{\textbf{H}}^{\mathbf{k}}=0$ at convergence. The gradient of the objective function $\Omega$ with respect to $\text{$\Gamma$}^{\mathbf{k}}$ can be computed efficiently, and the gradient with respect to ${\textbf{H}}^{\mathbf{k}}$ is obtained by symmetrizing the gradients with respect to $\text{$\Gamma$}^{\mathbf{k}}$. The detailed forms of the gradients are included below. All expressions are written using the Einstein summation convention. We define $G_{li}^{\textbf{k}}:=\frac{1-\delta_{li}}{\gamma_{i}^{\textbf{k}}-\gamma_{l}^{\textbf{k}}}$ where $\gamma_{i}^{\textbf{k}}$ is the $i$-th eigenvalue of $\mathbf{\text{$\Gamma$}}^{\mathbf{k}}$.
@@
To get the gradient with respect to ${\textbf{H}}^{\textbf{k}}$ from the gradient with respect to $\text{$\Gamma$}^{\textbf{k}}$, we can use the following:
$\frac{\partial\Omega}{\partial\text{Re}({\textbf{H}}^{\textbf{k}})}=\frac{\beta}{2}\left(\frac{\partial\Omega}{\partial\text{Re}\left(\text{$\Gamma$}^{\textbf{k}}\right)}+\left(\frac{\partial\Omega}{\partial\text{Re}\left(\text{$\Gamma$}^{\textbf{k}}\right)}\right)^{T}\right)$ (6)
$\frac{\partial\Omega}{\partial\text{Im}({\textbf{H}}^{\textbf{k}})}=\frac{\beta}{2}\left(\frac{\partial\Omega}{\partial\text{Im}\left(\text{$\Gamma$}^{\textbf{k}}\right)}-\left(\frac{\partial\Omega}{\partial\text{Im}\left(\text{$\Gamma$}^{\textbf{k}}\right)}\right)^{T}\right)$ (7)
With this method, we can converge systems at a wide range of chemical potentials. Our method has the same computational cost per iteration as standard electronic-structure convergence techniques, since the dominant cost remains the construction of the Fock matrix F, which is independent of the convergence scheme.
Note that this method can also be used to perform fixed-electron calculations as in canonical density functional theory by adding the gradient of the chemical potential $\mu$ with respect to the variational parameters and using the canonical free energy $E=E_{\text{DFT}}-TS_{\text{el}}$ as the objective function for optimization.16 This algorithm was also used to obtain the potential of zero charge (PZC) of metal slabs later in this work. The relevant gradients are as follows:
@@ II.2 Implicit Solvation Framework
Meaningful constant potential calculations in periodic systems must resolve one problem inherent to periodic boundary conditions. That is, the potential energy of charged unit cells in an infinite lattice diverges. In nearly all periodic electronic structure codes, the default way to handle this is by zeroing out the $G=0$ component of the electrostatic potential in reciprocal space, which is equivalent to introducing a uniform background charge that neutralizes the charge of the system.30, 29 It is known that the choice of neutralizing charge can affect qualitative predictions, so a physically informed choice should be used in calculations.51 As electrochemical reactions are performed in the presence of an electrolyte solution, the most logical way to enforce the net charge neutrality is by introducing ionic screening provided by electrolytes in solution. It is well-known that introducing linear ionic response (i.e., Debye screening) is already sufficient to neutralize the unit cell exactly.11
In this work, we have implemented a density-based polarizable continuum solvation model (PCM), which incorporates linear dielectric and ionic response. We demonstrate the implementation of two cavity functions. The first is a locally determined cavity based on the local linear PCM model (LPCM) in JDFTx and VASPsol.42, 9 The LPCM model is equivalent to computing the solvation potential via the linearized Poisson-Boltzmann equation. Secondly, we have implemented the charge-asymmetric nonlocally-determined local-electric (CANDLE) cavity function, known for its prediction of accurate energies for charged solutes, especially anions, which are systematically undersolvated in most PCM models.43 We provide more implementation details of these solvation models below.